# Uncertainty Propagation, and Parameter Calibration

### v1

@@ 4.2 Forward UQ with HMS
In this section, we demonstrate the use of the HMS for propagating uncertainty from the inputs to the regression rate as shown in Alg. 2. Approximating the distribution of the regression rate from the uncertain input space is useful for decision making and in this case to be able to bound the expected performance of the slab burner in terms of regression rate. We consider 3 areas of study for the slab burner since the regression rate is highly dependent on the location. Experimentally, the front part ($x\leq 40$mm) of the slab is expected to achieve the highest regression due to the ramp effects, the middle part ($40<x\leq 70$ mm) mostly achieves lower regression due to steady state combustion, and the rear part ($x>70$mm) also has higher regression due to re-circulation effects from the oxidizer flow. We do not expect the surrogate models to be able to match the experimental result however, as the 2D DNS does not either as shown previously in Fig. 7 where regression is low in the front part. In each forward propagation, we select boundary locations uniformly that adhere to the restrictions of the location. The velocity factor $V_{in}$ is also uniformly sampled and its range corresponds to an oxidizer flux of $G_{min}=5\frac{kg}{m^{s}}$ and $G_{max}=20\frac{kg}{m^{2}s}$, which are appropriate for this slab burner. The remaining two parameters are the latent heat of sublimation $l_{v}$ and the temperature exponent for reaction 249 ($b_{249}$) which are assumed normal with the mean being the nominal default value of the DNS and a variance that corresponds to a distribution that covers ±$\pm 25\%$ above or below the nominal value. The forward propagation uses traditional Monte-Carlo sampling and the results for the distributions of the regression rate as shown in Fig. 13. In the front of the slab, the estimate is a mean regression rate $\mu_{f}=0.022$ mm/s, in the middle of the slab $\mu_{m}=0.081$ mm/s, and in the rear $\mu_{b}=0.147$ mm/s. The front and rear distributions appear to have a more central tendency compared to the middle part where the density is more evenly distributed on a wider range of values.
Figure 10: $\mu_{f}=0.022,\sigma_{f}=0.021$mm/s
Figure 11: $\mu_{m}=0.081,\sigma_{m}=0.039$mm/s
Figure 12: $\mu_{b}=0.147,\sigma_{b}=0.029$mm/s
Figure 13: Forward propagation of uncertainty to the regression rate from uncertain inputs at 3 locations of the slab burner corresponding to $G=[5,20]\frac{kg}{m^{2}s}$. From left to right: $x_{slab}\leq 40$mm , $40<x_{slab}\leq 70$mm, $x_{slab}>70$mm. The 2D DNS estimates the regression to be lowest in the front and largest in the back of the slab. From experiments we know the front achieves the highest regression rate which is not captured by the DNS.
@@ 4.3 Bayesian model calibration using regression rate measurement data
In this section, we use the HMS surrogate model in conjunction with available experimental data (Fig. 6) to calibrate two key physical parameters of the DNS model: the latent heat of sublimation $l_{v}$ and the temperature exponent $b_{249}$. As described in Section 3.5, we apply Bayesian inference to calibrate these parameters while quantifying the associated uncertainty in model parameters. The prior distributions for the parameters are the same as used for the forward UQ, namely $b_{249}\sim\mathcal{N}(2.54,0.029)(-)$ and $l_{v}\sim\mathcal{N}(840890,4.91\times 10^{9})\frac{J}{kg}$. They are assumed normally distributed with the mean being the nominal default DNS value and a variance that results in the distribution covering values ±$\pm 25\%$ from that nominal value. The mean is taken from expert input and existing literature. To solve the Bayesian calibration problem, we employ the Random Walk Metropolis algorithm [87], a Markov Chain Monte Carlo (MCMC) method used to draw samples from the posterior distribution. To thoroughly explore the posterior distributions, we run 5 independent chains of 10,000 iterations each, starting from different initial points within the prior: both parameters at the edges of their feasible domain, both at their mean values, or one at the minimum and one at the maximum of their respective domains. The first 10% of each chain is discarded as “burn-in,” which is standard practice to allow the chain to converge towards the true posterior before sampling begins. Figs. 14, 15 present trace plots for one representative MCMC chain initialized with both parameters at the minimum of the prior range. Additionally, the autocorrelation plot (Fig. 17) is provided as a diagnostic tool to assess the independence of the samples. As shown, samples spaced 60 iterations apart exhibit minimal correlation, indicating that the MCMC chain is progressing efficiently and producing low-correlated, independent samples, and that the chain has appropriately explored the posterior distribution.
@@
To arrive at the final marginal posterior estimates (see Figs. 18, 19), we combine all 50000 samples from all 5 chains. A main conclusion from the results is that the default value for $l_{v}$ at the default value of 840890 J/kg is too large to allow the surrogate to achieve the experimental results in the front of the slab. Instead, the value should be lower ($\mu_{l_{v}}=410055\frac{J}{kg}$, $\sigma_{l_{v}}=205988\frac{J}{kg}$). The resulting posterior distribution for $b_{249}$ ($\mu_{b_{249}}=2.413$, $\sigma_{b_{249}}=0.177$) has a mean closer to the default value 2.54 from the DNS.
Figure 14: Bayesian calibration results: chain for $l_{v}$ starting with parameter at its prior minimum value.
Figure 15: Bayesian calibration results: chain for $b_{249}$ starting with parameter at its prior minimum value.
Figure 16: Bayesian calibration results: scatter plot of the chain for $l_{v}$ and $b_{249}$.
Figure 17: Bayesian calibration results: autocorrelation plot of the chain.
@@
Figure 18: Final marginal posterior distribution of $l_{v}$ using HMS and experimental data. We note that the posterior is far from the nominal value of 840890 $J/kg$ that the DNS uses as default. The calibration shows that $l_{v}$ should be lower ($\mu_{l_{v}}=410055\frac{J}{kg}$, $\sigma_{l_{v}}=205988\frac{J}{kg}$) for the HMS to get results closer to the data in the front part of the slab.
@@
Figure 19: Final marginal posterior distribution of $b_{249}$ using HMS and experimental data. The statistics of the distribution at ($\mu_{b_{249}}=2.413$, $\sigma_{b_{249}}=0.177$) are close to the default value of 2.54 used by the DNS.
To further demonstrate the impact of calibration, we show in Fig. 20, a forward UQ in two cases: “calibrated” and “not calibrated” for a specific oxidizer flux of $G=8.08\frac{kg}{m^{2}s}$, which is the same conditions used in calibration. The “not calibrated” case includes a forward UQ process with the same priors listed in Alg. 2. Instead, the “calibrated” case takes the prior distribution for $l_{v}$ and $b_{249}$ to be their posteriors from the Bayesian calibration. The result is a shift of the probability density for the regression rate towards higher values in all parts of the slab. The result verifies our previous observation: the DNS and subsequently the surrogates underestimate the regression rate in the front part of the slab, when used with the default values for the parameters in forward UQ. However, when we use calibrated priors for the parameters, we find the surrogate to produce samples for the front (now up to 0.14mm/s) that match the experiment. That also results in overestimating the regression rate in the middle and rear parts of the slab. The two calibrated parameters basically have to compensate for the model inadequacy between the 2D DNS and the experimental data and their effect is not sufficient to make the surrogate be simultaneously accurate in all slab locations. When uncalibrated, the surrogate underestimates the regression rate in the front part, correctly predicts the middle, and overestimates the rear part of the slab. When calibrated priors are used for the parameters, the surrogate correctly predicts the front, but further overestimates the middle and rear part of the slab. Overall, the conclusion is that there is not one set of ($l_{v},b_{249}$) that result in correctly predicting the regression rate in all locations of the slab. Calibration helps to identify the correct ranges to be used for specific locations. A next step would then be to have an accurate surrogate for different locations or conditions by simulating the DNS within the corresponding bounded parameter regimes.
Front slab
Middle slab
Rear slab
Figure 20: Forward UQ for a fixed velocity corresponding to $G=8.08\frac{kg}{m^{2}s}$ with default parameter priors and calibrated priors. The surrogate underestimates the regression rate in the front part of the slab, when used with the default values for the parameters in forward UQ. When we use calibrated priors for the parameters, we find the surrogate to produce samples for the front that match the experiment better.
@@ 5 Conclusion and Future Work
First, we generated an ensemble of 64 simulations that was informed by a series of sensitivity analyses to identify the most important reaction rate parameters. The simulation data was then used to develop two surrogate models: a Gaussian Process (GP) and a hyperparameter tuned Hierarchical Multiscale Surrogate (HMS). Our expectation was that the HMS would outperform the GP because of the multiscale effects of combustion. We tested both models under cross-validation for their ability to predict out of sample in parameter spaces they have not been trained on. We found that, overall, they performed very closely. The GP, on average, is expected to achieve an error of $1.59e^{-6}\%$ in the training set and $17.06\%$ when tested out of sample. The HMS, on average, is expected to achieve an error of $1.11\%$ in the training set and $16.62\%$ when tested out of sample. Our conclusion is that both of these surrogate models are acceptable for usage in UQ or otherwise. We do note however that the GP performance is sensitive to the choice of kernel and specifically kernels that cannot handle multiple scales (such as the rational quadratic) will result in poor models. HMS demonstrated a more stable behavior due to the usage of multiscale kernels and in some cases followed the simulation data closer. Neither model was able to fully capture nonlinearities. We justified this result from the fact that the surrogates are attempting prediction of a boundary quantity that is dependent on the combustion process, but none of their far field inputs they are trained on fully carry this multiscale information. The fact that both of these models are able to achieve an error $<20\%$ when predicting multiscale boundary quantities just from far field inputs is significant, particulary considering that they take a few $ms$ to run compared to 24 hours for the DNS.
@@
Second, we used the HMS in a forward UQ process to propagate uncertainty from the inputs to the QoI which is the regression rate. Assuming default priors for the input parameters (normal or uniform distributions as appropriate), we found that in the front of the slab, the estimate is a mean regression rate $\mu_{f}=0.022$ mm/s, in the middle of the slab $\mu_{m}=0.081$ mm/s, and in the rear $\mu_{b}=0.147$ mm/s. The front and rear distributions appear to have a more central tendency compared to the middle part where the density is more evenly distributed on a wider range of values. We also found that the surrogates (and by extension the 2D DNS) do not correctly predict the high regression rate in the front ramp of the slab burner.
Last, we used the Bayesian approach to calibrate two of the parameter inputs, the latent heat of sublimation $l_{v}$ and the temperature exponent $b_{249}$ with available experimental data. We found the posterior distribution of $l_{v}$ to have statistics $\mu_{l_{v}}=410055\frac{J}{kg}$, $\sigma_{l_{v}}=205988\frac{J}{kg}$, which correspond to values that are much lower than the default value. The resulting posterior distribution for $b_{249}$ ($\mu_{b_{249}}=2.413$, $\sigma_{b_{249}}=0.177$) includes similar values compared to the default value 2.54 from the DNS. We also conducted a study to compare the results of forward UQ when using the previous priors for the parameters and calibrated priors, for a fixed inlet velocity corresponding to the experiment conditions. We found that the two calibrated parameters have to compensate for the model inadequacy between the 2D DNS and the experimental data and their effect is not sufficient to make the surrogate be accurate in all slab locations simultaneously. When uncalibrated, the surrogate underestimates the regression rate in the front part, correctly predicts the middle, and overestimates the rear part of the slab. When calibrated priors are used for the parameters, the surrogate correctly predicts the front, but further overestimates the middle and rear part of the slab. Overall, the conclusion is that there is not one set of ($l_{v},b_{249}$) that result in correctly predicting the regression rate in all locations of the slab.
@@ Acknowledgements
The authors would also like to acknowledge the Tufts University High Performance Compute Cluster which was used for running the uncertainty analyses.
We thank Kolos Retfalvi, Jasper Stedman, Alex Post, and Joseph Cygan from the University at Buffalo for their help with acquiring and tracing the experimental data used for the Bayesian calibration.

### v2

@@ 3.2 Forward UQ with HMS
In this section, we demonstrate the use of the HMS for propagating uncertainty from the inputs to the regression rate as shown in Alg. 2. Approximating the distribution of the regression rate from the uncertain input space is useful for decision making and in this case to be able to bound the expected performance of the slab burner in terms of regression rate. We consider 3 areas of study for the slab burner since the regression rate is highly dependent on the location. Experimentally, the front part ($x\leq 30$mm) of the slab is expected to achieve the highest regression due to the ramp effects, the middle part ($30<x\leq 70$ mm) mostly achieves lower regression due to steady state combustion, and the rear part ($x>70$mm) also has higher regression due to re-circulation effects from the oxidizer flow. In each forward propagation, we select boundary locations uniformly that adhere to the restrictions of the location. The velocity factor $V_{in}$ is also uniformly sampled and its range corresponds to an oxidizer flux of $G_{min}=5\frac{kg}{m^{s}}$ and $G_{max}=20\frac{kg}{m^{2}s}$, which are appropriate for this slab burner. The remaining two parameters are the latent heat of sublimation $l_{v}$ and the activation energy for reaction 1 ($E_{1}$) which are assumed normal with the mean being the nominal default value of the DNS and a variance that corresponds to a distribution that covers ±$\pm 25\%$ above or below the nominal value. The forward propagation uses traditional Monte-Carlo sampling and the results for the distributions of the regression rate as shown in Figs. 11, 12, 13. In the front of the slab, the estimate is a mean regression rate $\mu_{f}=0.185$ mm/s, in the middle of the slab $\mu_{m}=0.093$ mm/s, and in the rear $\mu_{b}=0.142$ mm/s. All distributions appear to have a central tendency and follow the orders of magnitude we observe in the experiments at the corresponding locations.
Figure 11: Forward propagation of uncertainty to the regression rate from uncertain inputs at the front of the slab burner corresponding to $G=[5,20]\frac{kg}{m^{2}s}$.
Figure 12: Forward propagation of uncertainty to the regression rate from uncertain inputs at the middle of the slab burner corresponding to $G=[5,20]\frac{kg}{m^{2}s}$.
Figure 13: Forward propagation of uncertainty to the regression rate from uncertain inputs at the rear of the slab burner corresponding to $G=[5,20]\frac{kg}{m^{2}s}$.
@@ 3.3 Bayesian model calibration using regression rate measurement data
In this section, we use the HMS surrogate model in conjunction with available experimental data (Fig. 7) to calibrate two key physical parameters of the DNS model: the latent heat of sublimation $l_{v}$ and the activation energy $E_{1}$. As described in Section 2.5, we apply Bayesian inference to calibrate these parameters while quantifying the associated uncertainty in model parameters. The chosen prior distributions for the parameters $E_{1}\sim\mathcal{N}(15610,107.1)(cal/mol)$ and $l_{v}\sim\mathcal{N}(1e^{6},119369)\frac{J}{kg}$, which are wide enough to capture all possible values from literature [73]. To solve the Bayesian calibration problem, we employ the Random Walk Metropolis algorithm [74], a Markov Chain Monte Carlo (MCMC) method used to draw samples from the posterior distribution. To thoroughly explore the posterior distributions, we run 3 independent chains of 50,000 iterations each, starting from different initial points within the prior: both parameters at the minimum, mean, and maximum of their feasible domain. The first 20% of each chain is discarded as “burn-in,” which is standard practice to allow the chain to converge towards the true posterior before sampling begins. Figs. 14, 15 present trace plots for one representative MCMC chain initialized with both parameters at the mean of the prior range. Additionally, the autocorrelation plot (Fig. 16) is provided as a diagnostic tool to assess the independence of the samples. As shown, samples spaced 60 iterations apart exhibit minimal correlation, indicating that the MCMC chain is progressing efficiently and producing low-correlated, independent samples, and that the chain has appropriately explored the posterior distribution.
@@
To arrive at the final marginal posterior estimates (see Figs. 17, 18), we combine all 120000 samples from all 3 chains after burn-in. A main conclusion from the results is that the default values for both parameters are smaller than they should be to best allow the surrogate to approach the experimental results. Instead, the values should be higher based on the maximum aposteriori estimate (MAP) of the joint distribution ($l_{v,MAP}=1234973.57\frac{J}{kg},E_{1,MAP}=15847.24\frac{cal}{mol}$) as shown in Fig. 19.
Figure 14: MCMC chain for $E_{1}$ starting at prior mean value.
Figure 15: MCMC chain for $l_{v}$ starting at prior mean value.
Figure 16: Autocorrelation plot of $lv,E_{1}$ MCMC chains.
@@
Figure 17: Final marginal posterior distribution of $l_{v}$ after Bayesian calibration using HMS and experimental data. The calibration shows that $l_{v}$ should be much higher value (increased by $30\%+$) than the default value of the parameter in the DNS.
@@
Figure 18: Final marginal posterior distribution of $E_{1}$ after Bayesian calibration using HMS and experimental data. The calibration shows that $E_{1}$ should be slightly higher ( $1-2\%$) than the default value of the parameter in the DNS.
Figure 19: Final joint posterior distribution of $E_{1}$ and $l_{v}$ after Bayesian calibration using HMS and experimental data. The maximum aposteriori estimate (MAP) is at ($l_{v,MAP}=1234973.57\frac{J}{kg},E_{1,MAP}=15847.24\frac{cal}{mol}$)
@@ 4 Conclusion and Future Work
First, we generated an ensemble of 64 simulations that was informed by a series of sensitivity analyses to identify the most important reaction rate parameters. The simulation data was then used to develop two surrogate models: a Gaussian Process (GP) and a hyperparameter tuned Hierarchical Multiscale Surrogate (HMS). Our expectation was that the HMS would outperform the GP because of the multiscale effects of combustion. We tested both models under cross-validation for their ability to predict out of sample in parameter spaces they have not been trained on. We found that, overall, HMS is superior to GP for prediction. The GP, on average, is expected to achieve an error of $1.00e^{-8}\%$ in the training set and $28.03\%$ when tested out of sample. The HMS, on average, is expected to achieve an error of $0.63\%$ in the training set and $13.37\%$ when tested out of sample. Our conclusion is that HMS is superior to GP for usage in UQ or otherwise for this problem, due to its ability to better capture multiscale effects. HMS also demonstrated robust performance across the dataset. The fact that HMS is able to achieve an error $<15\%$ when predicting multiscale boundary quantities just from far field inputs is significant, particulary considering it takes a few $ms$ to run compared to 24 hours for the DNS.
@@
Second, we used the HMS in a forward UQ process to propagate uncertainty from the inputs to the QoI which is the regression rate. We assumed default priors for the input parameters (normal or uniform distributions as appropriate). At the front of the slab, the estimate is a mean regression rate $\mu_{f}=0.185$ mm/s, in the middle of the slab $\mu_{m}=0.093$ mm/s, and in the rear $\mu_{b}=0.142$ mm/s. All distributions appear to have a central tendency and follow the orders of magnitude we observe in the experiments at the corresponding locations.
Last, we used the Bayesian approach to calibrate two of the parameter inputs, the latent heat of sublimation $l_{v}$ and the activation energy $E_{1}$ with available experimental data. We found the maximum aposteriori estimate (MAP) for these two parameters to be at ($l_{v,MAP}=1234973.57\frac{J}{kg},E_{1,MAP}=15847.24\frac{cal}{mol}$). These MAP values are higher than the default used in the DNS (more than $30\%$ higher for $l_{v}$ and $1-2\%$ higher for $E_{1}$). Overall, the conclusion is that there is not one set of ($l_{v},E_{1}$) that result in correctly predicting the regression rate in all locations of the slab simultaneously.
@@ Acknowledgements
The authors would also like to acknowledge the Tufts University High Performance Compute Cluster which was used for running the uncertainty analyses. We thank Kolos Retfalvi, Jasper Stedman, Alex Post, and Joseph Cygan from the University at Buffalo for their help with acquiring and tracing the experimental data used for the Bayesian calibration.
Supplemental Material: Surrogate Comparison Figures
Figs. 20, 21 show additional comparisons between the GP and HMS for various simulation IDs across different folds.
Figure 20: Additional comparisons between GP and HMS for other simulation IDs from the 5-fold cross validation. The conclusions are the same as we outlined in Section 4 of the main paper.
Figure 21: Additional comparisons between GP and HMS for other simulation IDs from the 5-fold cross validation. The conclusions are the same as we outlined in Section 4 of the main paper.